Solve the following for the variable y:
a. 3y = 12
b.xyz=c
c. 3xy = G
d. c=ayb

Answer by checkley77(12844)

a. 3y = 12
Y=12/3=4 ANS.
b.xyz=c
Y=C/XZ ANS.
c. 3xy = G
Y=G/3X ANS.
d. c=ayb
Y=C/AB ANS.